OpenStudy (anonymous):

You are paid $7.45/hr at your part-time job. Your deductions are FICA (7.65%), federal tax withholding (10.65%), and state tax withholding (5.7%). Your travel expenses are $5.70 per day worked. You work 12 hours over a 3-day period of time. What is your net incomeafter travel expenses?

OpenStudy (janu16):

start off with the gross. 12 hrs @ 7.45/hr = 89.40 I will assume that you deduct your expenses before taxes. In that case, the taxable income is 89.40 - 3*5.70 = 72.30 Now for the tax bite. Total tax rate is .0765 + .1065 + .057 = .24 tax is thus .24*72.30 = 17.35 Subtract that, and your net income is 72.30-17.35 = 54.95 You could do it all in an equation, but it's easier to see by steps. In an algebraic expression, that would be (12*7.45-3*5.70)(1-(.0765+.1065+.057)) = 54.95

OpenStudy (janu16):

@commanderl17